# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ A. Multiply monogenic orders
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$, and $\mathcal{O}$ an arbitrary order of $K$, i.e., a subring of $\mathcal{O}_{K}$ with quotient field $K$. It follows from Corollary 5.3 above that up to $\mathbb{Z}$-equivalence, there are only finitely many $\alpha$ in $\mathcal{O}$ with $\mathcal{O}=\mathbb{Z}[\alpha]$. The order $\mathcal{O}$ is said to be $k$-times monogenic/precisely $k$ times monogenic/at most $k$ times monogenic if there are at least/precisely/at most $k$ pairwise $\mathbb{Z}$-inequivalent such generators $\alpha$ of $\mathcal{O}$ over $\mathbb{Z}$.
@@
For $n\geq 3$, Evertse and Győry (1985) proved that every order of a number field $K$ of degree $n$ is at most $(3\times 7^{2g})^{n-2}$-times monogenic, where $g$ denotes the degree of the normal closure of $K$. Note that $n\leq g\leq n!$. This was the first uniform bound of this type.
For fixed $n\geq 3$, we denote by $M(n)$ the smallest integer $k$ such that for every number field $K$ of degree $n$ and every order $\mathcal{O}$ of $K$, the order $\mathcal{O}$ is at most $k$ times monogenic. The above result of Evertse and Győry implies that for $n\geq 3$, the quantity $M(n)$ is finite and in fact, $M(n)\leq(3\times 7^{2n!})^{n-2}$.
The problem of estimating $M(3)$ can be reduced via index form equations to estimating the number of integer solutions of a Thue equation $|F(x,y)|=1$ with $F$ an integral cubic binary form. Bennett (2001) proved that such an equation has up to sign at most $10$ solutions. This gives the following.
Theorem A.1 (Bennett, 2001).
@@
Then, for $n\geq 4$, Evertse (2011) improved the bound of Evertse and Győry (1985).
Theorem A.2 (Evertse, 2011).
@@
The main tool in the proof of the general bound is the following result of Beukers and Schlickewei (1996): Let $a,b$ be non-zero complex numbers and $\Gamma$ a multiplicative subgroup of $\mathbb{C}^{*}$ of rank $r$. Then the equation $ax+by=1$ has at most $2^{16r+8}$ solutions in $x,y\in\Gamma$.
@@
Akhtari (2022) gave another, more direct proof for Theorem A.3, following the approach of Gaál, Pethő and Pohst (1996) (which in fact is going into the same direction as Bhargava’s approach but is less general), and combining this with her own upper bound for the number of solutions of quartic Thue equations.
Theorem A.2 is probably far from best possible in terms of $n$. We pose the following problem:
@@
Do there exist absolute constants $c_{1},c_{2}$ such that $M(n)<c_{1}n^{c_{2}}$ for all $n\geq 4$
@@
Theorem A.4 (Bérczes, Evertse and Győry, 2013).
@@
To see that the bound $3$ is optimal, let $K$ be a non-CM number field of degree $\geq 3$. Then the ring of integers of $K$ has infinitely many units $\varepsilon$ with $K=\mathbb{Q}(\varepsilon)$. For every of these $\varepsilon$ we obtain a two times monogenic order $\mathbb{Z}[\varepsilon]=\mathbb{Z}[\varepsilon^{-1}]$ of $K$.
Theorem A.4 is proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for these equations. So Theorems A.4 is ineffective, in the sense that its proof does not allow to determine the exceptional orders.
@@
Make Theorem A.4 effective.
@@
B. Generalizations to rationally monogenic orders
The results in Appendix A deal with monogenic orders, i.e., orders of the shape $\mathbb{Z}[\alpha]$ where $\alpha$ is an algebraic integer. There are generalizations of such orders, denoted $\mathbb{Z}_{\alpha}$, called rationally monogenic orders, attached to non-integral algebraic numbers $\alpha$. Further, the theorems stated in Appendix A have analogues for rationally monogenic orders. Before we can define rationally monogenic orders and state our theorems, we briefly go into some history and introduce the necessary terminology.
While in the results for monogenic orders, $\mathbb{Z}$-equivalence of algebraic integers plays an important role, for rationally monogenic orders we have to deal with $GL_{2}(\mathbb{Z})$-equivalence of algebraic numbers. Two algebraic numbers $\alpha,\beta$ are called $GL_{2}(\mathbb{Z})$-equivalent if $\beta=\frac{a\alpha+b}{c\alpha+d}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$. The correspondence with polynomials is as follows. Denote by $f_{\alpha}$ the primitive minimal polynomial of an algebraic number $\alpha$, i.e.,
(B.1) $f_{\alpha}=a_{0}X^{n}+\cdots+a_{n}=a_{0}(X-\alpha^{(1)})\cdots(X-\alpha^{(n)})\in\mathbb{Z}[X]$
where $a_{0}>0$, ${\rm gcd}(a_{0},\ldots,a_{n})=1$ and $\alpha^{(1)}=\alpha,,\ldots,\alpha^{(n)}$ are the conjugates of $\alpha$. Then if $\alpha,\beta$ are $GL_{2}(\mathbb{Z})$-equivalent then so are $f_{\alpha}$, $f_{\beta}$ while conversely, if $f_{\alpha},f_{\beta}$ are $GL_{2}(\mathbb{Z})$-equivalent, then $\alpha$ is $GL_{2}(\mathbb{Z})$-equivalent to a conjugate of $\beta$.
Let $\alpha$ be a non-zero, not necessarily integral algebraic number of degree $n\geq 3$, and $f_{\alpha}$ its primitive minimal polynomial given by (B.1). Define $\mathbb{Z}_{\alpha}$ to be the $\mathbb{Z}$-module with basis
This $\mathbb{Z}$-module was introduced by Birch and Merriman (1972), who observed that it is contained in the ring of integers of $\mathbb{Q}(\alpha)$, and that its discriminant is equal to the discriminant $D(f_{\alpha})$ of $f_{\alpha}$. Nakagawa (1989) showed that it is in fact an order of the field $\mathbb{Q}(\alpha)$, i.e., closed under multiplication. This order was further studied by Simon (2001, 2003) and Del Corso, Dvornicich and Simon (2005). As was very likely known at the time, the order $\mathbb{Z}_{\alpha}$ can be described more concisely as follows. Let $\mathcal{M}_{\alpha}$ be the $\mathbb{Z}$-module generated by $1,\alpha,\ldots,\alpha^{n-1}$. Then
$\mathbb{Z}_{\alpha}=\left\{\xi\in\mathbb{Q}(\alpha):\ \xi\mathcal{M}_{\alpha}\subseteq\mathcal{M}_{\alpha}\right\}.$
In the case that $\alpha$ is an algebraic integer of degree $n$, the powers $\alpha^{i}$ ($i\geq n$) belong to $\mathcal{M}_{\alpha}$, and thus, $\mathbb{Z}_{\alpha}=\mathcal{M}_{\alpha}=\mathbb{Z}[\alpha]$. Further, if $\alpha$, $\beta$ are non-zero $GL_{2}(\mathbb{Z})$-equivalent algebraic numbers, then $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$. Indeed, let $\beta=\frac{a\alpha+b}{c\alpha+d}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$. Then $\mathcal{M}_{\beta}=(c\alpha+d)^{1-n}\mathcal{M}_{\alpha}$ where $n=\deg\alpha$, and thus, $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$.
Nowadays, for a non-zero algebraic number $\alpha$ we call $\mathbb{Z}_{\alpha}$ the invariant order or invariant ring of $f_{\alpha}$.
From the effective result of Evertse and Győry (1991a) and the fact that $\mathbb{Z}_{\alpha}$ has the same discriminant as $f_{\alpha}$ the following analogue of Corollary 5.3 can be deduced:
Theorem B.1.
Let $\mathcal{O}$ be an order of a number field $K$ of degree $n$. Denote by $D_{\mathcal{O}}$ the discriminant of $\mathcal{O}$. Every $\alpha$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$ is $GL_{2}(\mathbb{Z})$-equivalent to some $\beta\in K$ of height $H(\beta)\leq c_{1}(n,|D_{\mathcal{O}}|)$, where $c_{1}(n,|D_{\mathcal{O}}|)$ is effectively computable in terms of $n$ and $|D_{\mathcal{O}}|$.
A consequence is that up to $GL_{2}(\mathbb{Z})$-equivalence there are only finitely many $\alpha\in K$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$, and that these can be effectively determined.
We call an order $\mathcal{O}$ of a number field $K$ rationally monogenic if there is $\alpha$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$. Clearly, monogenic orders are rationally monogenic. In Evertse (2023) it was shown that every number field $K$ of degree $\geq 3$ has infinitely many orders that are rationally monogenic but not monogenic. Simon (2001) gave various examples of number fields of degree $\geq 4$ that are not rationally monogenic, i.e., whose rings of integers are not rationally monogenic.

### v2

@@ 10.1. Monogenic orders
In this subsection, we consider (10.1). Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$, and $\mathcal{O}$ an arbitrary order of $K$, i.e., a subring of $\mathcal{O}_{K}$ with quotient field $K$. It follows from Theorem 5.6 above (in an effective form) that up to $\mathbb{Z}$-equivalence, there are only finitely many $\alpha\in\mathcal{O}$ with $\mathcal{O}=\mathbb{Z}[\alpha]$. The order $\mathcal{O}$ is said to be $k$-times monogenic/precisely $k$ times monogenic/at most $k$ times monogenic if there are at least/precisely/at most $k$ pairwise $\mathbb{Z}$-inequivalent such generators $\alpha$ of $\mathcal{O}$ over $\mathbb{Z}$.
@@
For fixed $n\geq 3$, we denote by $M(n)$ the smallest integer $k$ such that for every number field $K$ of degree $n$ and every order $\mathcal{O}$ of $K$, the order $\mathcal{O}$ is at most $k$ times monogenic. We start with recalling an old result of ours.
Theorem 10.1 (Evertse and Győry, 1985).
Let $K$ be a number field of degree $n\geq 3$, and suppose that its normal closure has degree $g$. Then every order of $K$ is at most $(3\times 7^{2g})^{n-2}$ times monogenic.
In particular, $M(n)$ is finite, and $M(n)\leq(3\times 7^{2n!})^{n-2}$.
This was deduced from an upper bound for the number of solutions of $S$-unit equations, obtained shortly before by the first author, see Evertse (1984a).
There are now much better upper bounds for $M(n)$. The problem of estimating $M(3)$ can be reduced via index form equations to estimating the number of integer solutions of a Thue equation $|F(x,y)|=1$ with $F$ an integral cubic binary form. Bennett (2001) proved that such an equation has up to sign at most $10$ solutions. This gives the following.
Theorem 10.2 (Bennett, 2001).
@@
For $n\geq 4$, the first author improved the bound of Theorem 10.1 as follows.
Theorem 10.3 (Evertse, 2011).
@@
The main tool in the proof is an important improvement and generalization of the first author’s result from 1984, due to Beukers and Schlickewei (1996), see Theorem 10.6 in Section 10.2.
@@
Akhtari (2022) gave another, more direct proof for Theorem 10.4, following the approach of Gaál, Pethő and Pohst (1996) (which in fact is going into the same direction as Bhargava’s approach but is less general), and combining this with her own upper bound for the number of solutions of quartic Thue equations.
Theorem 10.3 is probably far from best possible in terms of $n$. We pose the following problem:
@@
Do there exist absolute constants $c_{1},c_{2}$ such that $M(n)<c_{1}n^{c_{2}}$ for all $n\geq 4$ ?
The best lower bound we could find is due to Miller-Sims and Robertson (2005). Let $p$ be a prime number, $\zeta_{p}$ a primitive $p$-th root of unity, and $K_{p}$ the associated real cyclotomic field, i.e., $\mathbb{Q}(\zeta_{p}+\zeta_{p}^{-1})$. Then $K_{p}$ has degree (p1)/2(p-1)/2, and its ring of integers is $\mathcal{O}_{p}:=\mathbb{Z}[\zeta_{p}+\zeta_{p}^{-1}]$. They proved that if $p\geq 7$ then $\mathbb{Z}[\alpha]=\mathcal{O}_{p}$ is satisfied by $\alpha=\zeta_{p}^{k}+\zeta_{p}^{-k}$, $(\zeta_{p}^{k}+\zeta_{p}^{-k}+b)^{-1}$ (b=1,0,1,2b=-1,0,1,2, $k=1,\ldots,(p-1)/2$). If $p=7$ then among these numbers there are precisely nine pairwise $\mathbb{Z}$-inequivalent ones and by a result of Gaál and Schulte (1989) these are up to $\mathbb{Z}$-equivalence the only numbers $\alpha$ with $\mathbb{Z}[\alpha]=\mathcal{O}_{7}$. If $p\geq 11$ then all these numbers are pairwise $\mathbb{Z}$-inequivalent and thus, the order $\mathcal{O}_{p}$ is $5(p-1)/2=5[K_{p}:\mathbb{Q}]$ times monogenic.
@@
Theorem 10.5 (Bérczes, Evertse and Győry, 2013).
@@
To see that this is optimal, let $K$ be a non-CM number field of degree $\geq 3$. Then the ring of integers of $K$ has infinitely many units $\varepsilon$ with $K=\mathbb{Q}(\varepsilon)$. For every of these $\varepsilon$ we obtain a two times monogenic order $\mathbb{Z}[\varepsilon]=\mathbb{Z}[\varepsilon^{-1}]$ of $K$.
Theorem 10.5 is proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for these equations. So Theorem 10.5 is ineffective, in the sense that its proof does not allow to determine the exceptional orders.
@@
Make Theorem 10.5 effective.
@@
10.2. Outlines of the proofs of Theorems 10.3 and 10.5
a
We start with recalling some auxiliary results from the literature.
Theorem 10.6 (Beukers and Schlickewei, 1996).
Let $F$ be a field of characteristic 0, and $\Gamma$ a multiplicative subgroup of $F^{*}\times F^{*}$ of rank $r$. Then the equation $x+y=1$ has at most $2^{8r+8}$ solutions $(x,y)\in\Gamma$.
Corollary 10.7.
Let $F$ be a field of characteristic 0, let $m\geq 1$, and let $\Gamma$ be a multiplicative subgroup of $(F^{*})^{2m}$ of rank $r$. Then there are at most $2^{8(r+2m-1)}$ tuples $(x_{1},y_{1},\ldots,x_{m},y_{m})\in\Gamma$ satisfying
(10.3) $x_{i}+y_{i}=1\ \ \text{for }i=1,\ldots,m.$
This result is easily deduced from Theorem 10.6 using induction on $m$, see Evertse (2011), or Evertse and Győry (2017), Corollary 4.3.5.
Theorem 10.8.
Let $F$ be a field of characteristic 0, let $m\geq 1$, and let $\Gamma$ be a multiplicative subgroup of $(F^{*})^{m}$. Then there are at most finitely many tuples $(x_{1},\ldots,x_{m})\in\Gamma$ satisfying
(10.4) $\left\{\begin{array}[]{l}x_{1}+\cdots+x_{m}=1,\\ x_{i_{1}}+\cdots+x_{i_{t}}\not=0\ \text{for each non-empty subset $\{i_{1},\ldots,i_{t}\}$ of $\{1,\ldots,m\}$.}\end{array}\right$.
This was proved by Evertse (1984b) and van der Poorten and Schlickewei (1982, 1991), combining Schmidt’s and Schlickewei’s Subspace Theorem from Diophantine approximation with a specialization argument. We note that Theorem 10.8 is ineffective, hence so are its consequences. Although we will not need these here, we mention that there are explicit upper bounds for the number of solutions of (10.4) depending only on $m$ and on $r:={\rm rank}\,\Gamma$, see Evertse, Schlickewei and Schmidt (2002) or Amoroso and Viada (2009), who obtained the up to now best upper bound $(8m)^{4m^{4}(m+r+1)}$.
Theorem 10.9.
Let $F$ be a field of characteristic 0, and $\Gamma$ a multiplicative subgroup of $F^{*}\times F^{*}$. Then there are only finitely many pairs $(a,b)\in F^{*}\times F^{*}$ such that $a+b=1$, and such that $ax+by=1$ has three solutions $(x,y)\in\Gamma$, the pair $(1,1)$ included.
Idea of proof.
This is basically a result of Evertse, Győry, Stewart, and Tijdeman (1988), see also Evertse and Győry (2015), Theorem 6.1.6. The idea is as follows. Suppose that there are $(x_{1},y_{1}),(x_{2},y_{2})\in\Gamma$, distinct from each other and distinct from $(1,1)$, such that $ax_{i}+by_{i}=1$ for $i=1,2$. Then
$\left|\begin{array}[]{lll}1&1&1\\ 1&x_{1}&y_{1}\\ 1&x_{2}&y_{2}\end{array}\right|=0.$
Expand the determinant, divide by a term to obtain a five term sum equal to $1$, consider all possible partitions into minimal vanishing subsums, and apply Theorem 10.8 to each of them. ∎
Let $K$ be a number field of degree $n\geq 3$. Denote by $x\mapsto x^{(i)}$ ($i=1,\ldots,n$) the embeddings of $K$ in $G$, where $G$ is the normal closure of $K$. For $\alpha$ with $\mathbb{Q}(\alpha)=K$ and $i=3,\ldots,n$, define
$x_{i}(\alpha)=\frac{\alpha^{(i)}-\alpha^{(1)}}{\alpha^{(2)}-\alpha^{(1)}},\ \ y_{i}(\alpha)=\frac{\alpha^{(2)}-\alpha^{(i)}}{\alpha^{(2)}-\alpha^{(1)}}$
and the tuple
$\kappa(\alpha):=(x_{3}(\alpha),y_{3}(\alpha),\ldots,x_{n}(\alpha),y_{n}(\alpha)).$
In addition, we need a few simple lemmas. We call $\alpha,\beta$ $\mathbb{Q}$-equivalent if $\beta=\lambda\alpha+a$ for some $\lambda\in\mathbb{Q}^{*}$, $a\in\mathbb{Q}$.
Lemma 10.10.
Let $\alpha,\beta$ with $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$.
(i) $\kappa(\alpha)=\kappa(\beta)$ $\Longleftrightarrow$ $\alpha,\beta$ are $\mathbb{Q}$-equivalent.
(ii) Assume in addition that $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$ and that $\alpha,\beta$ are $\mathbb{Q}$-equivalent. Then $\alpha,\beta$ are $\mathbb{Z}$-equivalent.
Proof.
(i) Clearly, $\kappa(\alpha)=\kappa(\beta)$ if and only if $(\alpha^{(i)},\beta^{(i)})$ ($i=1,\ldots,n$) are collinear, i.e., $\beta^{(i)}=\lambda\alpha^{(i)}+a$ ($i=1,\ldots,n$) for some $\lambda\in G^{*}$, $a\in G$. One easily shows that this is posible only if $\lambda,a$ are invariant under Galois action, i.e., lie in $\mathbb{Q}$.
(ii) Our assumption $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$ implies that $\beta=f(\alpha)$ for some unique polynomial $f\in\mathbb{Q}[X]$ of degree <n<n, and then $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$ implies $f\in\mathbb{Z}[X]$. So if $\alpha,\beta$ are $\mathbb{Q}$-equivalent, then $\beta=\lambda\alpha+a$ with $\lambda,a\in\mathbb{Z}$. By interchanging the role of $\alpha,\beta$ we see that $\lambda^{-1}\in\mathbb{Z}$, hence $\lambda=\pm 1$. ∎
Lemma 10.11.
Let $\alpha,\beta$ be such that $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$ and $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$. Then
$\frac{\beta^{(i)}-\beta^{(j)}}{\alpha^{(i)}-\alpha^{(j)}}\in\mathcal{O}_{G}^{*}\ \text{for }i,j=1,\ldots,n,\,i\not=j.$
Proof.
Use $\beta=f(\alpha)$, $\alpha=g(\beta)$ for some $f,g\in\mathbb{Z}[X]$. ∎
Sketch of the proof of Theorem 10.3.
Let $\mathcal{O}$ be an order of $K$. Note that for $\alpha\in K$ with $K=\mathbb{Q}(\alpha)$ we have relations
$x_{i}(\alpha)+y_{i}(\alpha)=1\ \ (i=3,\ldots,n)$
where
$x_{i}(\alpha)=\frac{\alpha^{(i)}-\alpha^{(1)}}{\alpha^{(2)}-\alpha^{(1)}},\ \ y_{i}(\alpha)=\frac{\alpha^{(2)}-\alpha^{(i)}}{\alpha^{(2)}-\alpha^{(1)}}.$
It was proved in Evertse (2011), see also Evertse and Győry (2017), pages 206–208, that if one restricts to $\alpha$ with $\mathbb{Z}[\alpha]=\mathcal{O}$, then the set of tuples
$\{\kappa(\alpha):\,\mathbb{Z}[\alpha]=\mathcal{O}\}$
generates a multiplicative subgroup of $(G^{*})^{2n-4}$ of rank at most n(n1)/2n(n-1)/2. In the deduction of this we used a refinement of Lemma 10.11. Now an application of Corollary 10.7 and Lemma 10.10 implies Theorem 10.3. ∎
In the proof of Theorem 10.5 we need the following lemma. Call $\alpha_{1}$ $k$-special if $\alpha_{1}\in O_{K}$, $K=\mathbb{Q}(\alpha_{1})$ and there are $\alpha_{2},\ldots,\alpha_{k}$ such that $\alpha_{1},\ldots,\alpha_{k}$ are pairwise $\mathbb{Z}$-inequivalent and $\mathbb{Z}[\alpha_{1}]=\mathbb{Z}[\alpha_{2}]=\cdots=\mathbb{Z}[\alpha_{k}]$.
Lemma 10.12.
Let $\mathcal{C}$ be a $\mathbb{Q}$-equivalence class of $2$-special numbers. Then $\mathcal{C}$ is the union of finitely many $\mathbb{Z}$-equivalence classes.
Proof.
For the somewhat involved argument we refer to Bérczes, Evertse and Győry (2013) or Evertse and Győry (2017), Lemma 9.5.6. ∎
Sketch of the proof of Theorem 10.5.
We have to prove that there are only finitely many orders $\mathbb{Z}[\alpha]$ such that $\alpha$ is $3$-special. It suffices to show that the $3$-special $\alpha$ lie in finitely many $\mathbb{Z}$-equivalence classes. We sketch the argument.
Let $\alpha\in O_{K}$ be $3$-special. Pick $\beta,\gamma$ such that $\alpha,\beta,\gamma$ are pairwise $\mathbb{Z}$-inequivalent, and $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]=\mathbb{Z}[\gamma]$. For any three distinct indices $i,j,k$ from $\{1,\ldots,n\}$, define
$\varepsilon_{ijk}=\frac{(\beta^{(i)}-\beta^{(j)})/(\beta^{(i)}-\beta^{(k)})}{(\alpha^{(i)}-\alpha^{(j)})/(\alpha^{(i)}-\alpha^{(k)})},\ \ \eta_{ijk}=\frac{(\gamma^{(i)}-\gamma^{(j)})/(\gamma^{(i)}-\gamma^{(k)})}{(\alpha^{(i)}-\alpha^{(j)})/(\alpha^{(i)}-\alpha^{(k)})}.$
Then by Lemma 10.11, the equation
$\frac{\alpha^{(i)}-\alpha^{(j)}}{\alpha^{(i)}-\alpha^{(k)}}x+\frac{\alpha^{(j)}-\alpha^{(k)}}{\alpha^{(i)}-\alpha^{(k)}}y=1\ \ \text{in }x,y\in\mathcal{O}_{G}^{*}$
has three solutions
$(1,1),\ (\varepsilon_{ijk},\varepsilon_{kji}),\ (\eta_{ijk},\eta_{kji}).$
If for all $i,j,k$ and all $\alpha,\beta,\gamma$ as above these three solutions were distinct, we could conclude from Theorem 10.9 that there is a finite set $S\SS$ such that $S\alpha^{(i,j,k)}\in\SS$ for all $i,j,k$ and all $3$-special $\alpha$. It need not be true, however, that in all cases these three solutions are distinct. However, by means of a combinatorial argument, worked out in Bérczes, Evertse and Győry (2013) or Evertse and Győry (2017), pp. 211–216 we deduce that the existence of a finite set $S\SS$ as above still holds. Now Lemma 10.10 (i) implies that the $3$-special numbers $\alpha$ lie in only finitely many $\mathbb{Q}$-equivalence classes. Finally, Lemma 10.12 implies that the $3$-special numbers lie in only finitely many $\mathbb{Z}$-equivalence classes. ∎
10.3. Generalizations for rationally monogenic orders
a
The theorems stated in Subsection 10.1 have analogues for rationally monogenic orders. For the necessary terminology and properties we refer to Subsection 5.6.
For a not necessarily integral algebraic number $\alpha$ of degree $n\geq 3$ we define
$\mathcal{M}_{\alpha}$ $:=\{x_{0}+x_{1}\alpha+\cdots+x_{n-1}\alpha^{n-1}:\,x_{0},\ldots,x_{n-1}\in\mathbb{Z}\},$
$\mathbb{Z}_{\alpha}$ $:=\{\xi\in\mathbb{Q}(\alpha):\,\xi\mathcal{M}_{\alpha}\subseteq\mathcal{M}_{\alpha}\}.$
Recall that $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$ if $\alpha$ and $\beta$ are $GL_{2}(\mathbb{Z})$-equivalent.
An order $\mathcal{O}$ of a number field $K$ is called rationally monogenic if $\mathcal{O}=\mathbb{Z}_{\alpha}$ for some algebraic number $\alpha$. As observed in Subsection 5.6, if $\alpha$ is an algebraic integer, then $\mathbb{Z}_{\alpha}=\mathbb{Z}[\alpha]$. Thus, monogenic orders are rationally monogenic. We further recall that rationally monogenic orders are primitive, i.e., they cannot be expressed as $\mathbb{Z}+a\mathcal{O}^{\prime}$ for some integer $a>1$ and order $\mathcal{O}^{\prime}$.